# Optimal Stopping with a Predicted Prior

### v1

@@ 6 Algorithm for MaxExp: Proof of Theorem 1
By the implicit sharding method (Theorem 7), it suffices to design an algorithm $\mathcal{A}(\tilde{F},\theta)$ under the bi-criteria framework, and prove the desired consistency and robustness guarantees as the number of values $n$ tends to infinity. Further, by Corollary 9, the robustness guarantee holds for any robust threshold $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ where $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$.
The rest of the section focuses on the consistency guarantee, for any fixed robustness parameter $\beta\in[0,\frac{1}{e}]$ and the corresponding $\lambda_{1},\lambda_{2}$. Since the predicted prior is correct, i.e., $\tilde{F}=F$, we will only use $F$ throughout the section. To formally study the limit as $n\to\infty$, we will design the threshold function for the extremal distribution $F^{n}$. More precisely, we will design a threshold function $\theta$ and analyze the consistency of algorithm $\mathcal{A}(F,\theta^{\nicefrac{{1}}{{n}}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ for $n$ values, and in particular, the limit when $n$ tends to infinity.
As a standard approach, we design the algorithm so that for any $\ell\geq 0$, it holds
$\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]~\geq~\alpha\cdot\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Opt}}\geq\ell\,\big]~=~\alpha\cdot\big(1-F(\ell)^{n}\big)~.$ (3)
This implies the consistency guarantee as
@@ 6.1 Sufficient Conditions for Infinitely Many Values
We consider a general threshold function $\theta$ and derive an expression for the left-hand side of Equation 3, i.e., the probability that the algorithm accepts a value that is at least $\ell$.
@@
For any $y\in[0,1]$, if we let $n$ tend to infinity while keeping the same $F(\ell)^{n}=y$, the desired Inequality (3) becomes
$\int_{y}^{1}\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),q\big\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t\,\mathrm{d}q~\geq~\alpha\cdot(1-y)~,$
where we use the fact that $\big(1-t+tz^{\nicefrac{{1}}{{n}}}\big)^{n-1}\to z^{t}$ as $n$ approaches infinity. [^7] Rigorously, suppose we have $\alpha^{*}$-consistency in the limit when $n$ tends to infinity and consider any $\varepsilon>0$. By the uniform convergence of the stated limit for $q\in[\varepsilon,1]$, we can apply implicit sharding to ensure Equation 3 for $\alpha=\alpha^{*}(1-\varepsilon)$ and any $q=F(\ell)^{n}\in[\varepsilon,1]$. Finally, Equation 3 holds for $\alpha^{*}(1-\varepsilon)^{2}$ when $q=F(\ell)^{n}\in[0,\varepsilon]$ because the left-hand-side is decreasing in $q$, while the right-hand-side differs by at most a $1-\varepsilon$ factor.
@@
$g(q)~\coloneqq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),q\big\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~.$
Then, it suffices to design a threshold function $\theta$ such that for any $q\in[0,1]$, it holds
@@
6.2 Proof of Theorem 1
Next, we focus on a robust threshold function $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$, and for ease of notation we continue to write $\theta$ instead of $\theta\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$, restricting $\theta$ from now on such that
@@ 6.2 Proof of Theorem 1
    $\theta(t)=1$ for $t\in[0,\lambda_{1})$; and
@@
    $\theta(t)=0$ for $t\in(\lambda_{2},1]$.
Further, recall the assumed properties in Theorem 1:
  1. 3)
    $\theta$ is left-continuous, which implies $\theta(\lambda_{1})=1$, and non-increasing; and
  2. 4)
@@
    $\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\cdot\theta(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t~\geq~\alpha\cdot\theta(z)~.$
Consider any $q$ with $\theta(\lambda_{2})\leq q\leq\theta(\lambda_{1})=1$. Let $z=\theta^{-1}(q)$, and then by property 4), we have
$\alpha$ $~\leq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\theta(\max\{s,\theta^{-1}(q)\})^{t}}{\theta(\theta^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t$
$~=~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\theta(s),\theta(\theta^{-1}(q))\big\}^{t}}{\theta(\theta^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t.$
By left-continuity of $\theta$ and the definition of $\theta^{-1}$, we have $\theta(\theta^{-1}(q))\geq q$. Therefore,
$\alpha~\leq~\int_{\theta^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\{\theta(s),q\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~=~g(q).$
Finally, we verify $g(q)\geq\alpha$ for $q\in[0,\theta(\lambda_{2}))$ by reducing it to the first case.
@@
For any $q\in[0,\theta(\lambda_{2})]$, we have
$g(q)\geq g\big(\theta(\lambda_{2})\big).$
@@
For any $q$ in the range $0,\theta(\lambda_{2})]$, we have $\theta^{-1}(q)=\lambda_{2}$. Also, observe that $\theta(s)\geq\theta(\lambda_{2})\geq q$ for $s\in[0,\lambda_{2}]$, and $\theta(s)=0$ for $s\in[\lambda_{2},1]$. Hence,
@@
which is decreasing in $q$. ∎
@@ 7 Algorithm for MaxProb: Proof of Theorem 2
We first derive an explicit expression for the probability that algorithm $\mathcal{A}(\tilde{F},\theta)$ accepts the maximum value when the predicted prior is correct.
@@
For any $n\in\mathbb{N}^{+}$, any prior $F$, and any threshold function $\theta$, algorithm $\mathcal{A}(\tilde{F}=F,\theta)$ is
$\Gamma_{n}(\theta)\coloneq\int_{0}^{1}\bigg(\int_{s}^{1}\frac{(1-t+t\theta(s))^{n}-t\theta(s)^{n}}{t(1-t)}\,\mathrm{d}t\;-\;\theta(s)^{n}\bigg)\,\mathrm{d}s$
competitive for MaxProb.
@@
Changing the order of integration and simplifying yields the stated expression for $\Gamma_{n}(\theta)$. ∎
@@
For any $n\in\mathbb{N}^{+}$ and any $\lambda_{1},\lambda_{2}$ with $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$, define threshold function $\theta_{n}^{*}$ such that for any $s\in[0,1]$, $\theta_{n}^{*}(s)$ is the unique solution to
$\int_{s}^{1}\frac{\big(1-t+t\theta_{n}^{*}(s)\big)^{n-1}-\theta_{n}^{*}(s)^{n-1}}{1-t}\,\mathrm{d}t~=~\theta_{n}^{*}(s)^{n-1}~.$ (4)
We have
$\Gamma_{n}\big(\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~\geq~-\lambda_{1}\ln\lambda_{1}+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~,$
where equality holds when $n$ tends to infinity.
Note that $\theta_{n}^{*}$ is exactly the threshold function used in the continuous-time Gilbert-Mosteller algorithm (Example 2). We defer the proofs of some lemmas to Appendix A.
@@
By the expression for $\Gamma_{n}(\theta)$ in Lemma 13, we derive the definition of $\theta_{n}^{*}$ from the first-order condition for $\theta(t)$, i.e.,
[…]
Further, $\theta_{n}^{*}(s)$ is decreasing in $s$ because Equation 4 can be rewritten as
$\int_{s}^{1}\frac{\big(\tfrac{1-t}{\theta_{n}^{*}(s)}+t\big)^{n-1}-1}{1-t}\,\mathrm{d}t~=~1~,$ (5)
except for the boundary condition $\theta_{n}^{*}(1)=0$.
In sum, $\theta_{n}^{*}$ is a threshold function satisfying the first-order conditions for $t\in[0,1]$. Therefore, $\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ is a threshold function satisfying the first-order conditions for $t\in[\lambda_{1},\lambda_{2}]$ on one hand, and the restrictions of being $1$ for $t\in[0,\lambda_{1})$ and 0 for $t\in(\lambda_{2},1]$ on the other hand.
We next write the competitive ratio $\Gamma_{n}\big(\theta_{n}^{*}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ as
$\int_{0}^{\lambda_{1}}\bigg(\int_{s}^{1}\frac{1}{t}\,\mathrm{d}t-1\bigg)\,\mathrm{d}s+\int_{\lambda_{1}}^{\lambda_{2}}\bigg(\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n}-t\theta_{n}^{*}(s)^{n}}{t(1-t)}\,\mathrm{d}t-\theta_{n}^{*}(s)^{n}\bigg)\,\mathrm{d}s+\int_{\lambda_{2}}^{1}\frac{(1-t)^{n-1}}{t}\,\mathrm{d}s~.$
We now focus on the limit case when $n$ tends to infinity. This is without loss of generality, either by the implicit sharding technique, or by the fact that the above ratio is non-increasing in $n$ (see Lemma 19 in Section A.1). In particular, the latter means that the stated consistency and robustness guarantees also hold directly for the bi-criteria algorithm without implicit sharding.
When $n\to\infty$, the first term equals $\lambda_{1}\ln\lambda_{1}$ by basic calculus, and the last term tends to 0. It remains to show that, as $n\to\infty$, the integrand of the second term converges to
@@
By the definition of $\theta_{n}^{*}$ (Equation 4), the integrand equals
$\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n}-t\theta_{n}^{*}(s)^{n}}{t(1-t)}\,\mathrm{d}t\,-\,\theta_{n}^{*}(s)\int_{s}^{1}\frac{\big(1-t+t\theta_{n}^{*}(s)\big)^{n-1}-\theta_{n}^{*}(s)^{n-1}}{1-t}\,\mathrm{d}t$
$\qquad={}\int_{s}^{1}\frac{(1-t+t\theta_{n}^{*}(s))^{n-1}}{t}\,\mathrm{d}t~.$
Finally, we prove in Section A.2 (Lemma 20) that
@@
$\theta_{n}^{*}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\!\left(\frac{1}{n}\right)$

### v2

@@ 6 Algorithm for MaxExp: Proof of Theorem 1
In this section we prove Theorem 1. We fix a $\beta\in[0,\nicefrac{{1}}{{e}}]$, and fix $0\leq\lambda_{1}\leq\lambda_{2}\leq 1$ as the two roots of $\lambda\ln\lambda=\beta$. Once there exists a non-increasing function $\bar{\theta}\colon[\lambda_{1},\lambda_{2}]\to[0,1]$ that satisfies Equation (1), with a little abuse of notation, we extend the definition of $\bar{\theta}$ to the entire interval $[0,1]$ by setting $\bar{\theta}(t)=1$ for $t<\lambda_{1}$ and $\bar{\theta}(t)=0$ for $t>\lambda_{2}$.
By Theorem 9, it suffices to construct an algorithm that is both $\alpha$-consistent and $\beta$-robust as the number of values $n$ tends to infinity. We show that $\mathcal{A}(\tilde{F},\bar{\theta}^{\nicefrac{{1}}{{n}}})$ satisfies these guarantees. The $\beta$-robustness directly follows from Lemma 7, since $\bar{\theta}^{\nicefrac{{1}}{{n}}}=\bar{\theta}^{\nicefrac{{1}}{{n}}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ by the extension above.
The rest of the section focuses on the consistency guarantee. Since the predicted prior is correct, i.e., $\tilde{F}=F$, we will write it as $F$ throughout the section. The remaining task is to show that $\mathcal{A}(F,\bar{\theta}^{\nicefrac{{1}}{{n}}})$ is $\alpha$-competitive as $n\to\infty$.
As a standard approach in competitive analysis, we show that for any $\ell\geq 0$, it holds
$\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Alg}}\geq\ell\,\big]~\geq~\alpha\cdot\operatorname{\mathbf{Pr}}\big[\,x_{\textsc{Opt}}\geq\ell\,\big]~=~\alpha\cdot\big(1-F(\ell)^{n}\big)~,$ (4)
i.e., $x_{\textsc{Alg}}$ stochastically dominates $\alpha\cdot x_{\textsc{Opt}}$. This implies the desired consistency
@@
Remark 1.
The algorithm designed in this section does not use the theoretically optimal threshold $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$ derived in Section 3.3; that threshold maximizes consistency among all algorithms with robust thresholds of the form $\theta|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}$. For the MaxExp objective, $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ is given by $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}(t)=F(x^{*}(t))$, where $x^{*}(t)$ is the unique solution of the Bellman equation (see Equation 17), so $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ depends on the prior distribution $F$ through quantities such as the conditional expectation $\operatorname{\mathbf{E}}_{y\sim F}[y\mid y\geq x]$ and the quantile $F(x)$. As a result, the consistency guarantees $\operatorname{\mathbf{Pr}}[x_{\textsc{Alg}}\geq\ell]$ under $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ are expressed in terms of these prior-dependent objects. In this section we instead work with a suboptimal but analytically more tractable threshold that is independent of the prior $F$, and we analyze its consistency relative to $\theta_{n}^{\scriptscriptstyle\mathrm{exp}}$ for each fixed prior.
The proof consists of two parts. First, we simplify Equation 4 to give a sufficient condition that $\bar{\theta}$ needs to satisfy. Then, we verify that $\bar{\theta}$ satisfies these conditions.
@@ 6.1 Sufficient Conditions for Infinitely Many Values
First, to simplify the left-hand side of Equation 4, we use the following lemma to calculate the probability a bi-criteria algorithm $\mathcal{A}(F,\theta^{\nicefrac{{1}}{{n}}})$ accepts a value that is at least $\ell$.
@@
Now, focus on algorithm $\mathcal{A}(F,\bar{\theta}^{\nicefrac{{1}}{{n}}})$. For any $y\in[0,1]$, if we let $n\to\infty$ while keeping the same $F(\ell)^{n}=y$, the desired Inequality (4) becomes
$\int_{y}^{1}\int_{\bar{\theta}^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\{\bar{\theta}(s),q\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t\,\mathrm{d}q~\geq~\alpha\cdot(1-y)~,$
where we use the fact that $\big(1-t+tz^{\frac{1}{n}}\big)^{n-1}\to z^{t}$ as $n$ approaches infinity.  [^8] Rigorously, suppose we have $\alpha^{*}$-consistency in the limit when $n\to\infty$ and consider any $\varepsilon>0$. By the uniform convergence of the stated limit for $q\in[\varepsilon,1]$, we can apply implicit sharding to ensure Equation 4 for $\alpha=\alpha^{*}(1-\varepsilon)$ and any $q=F(\ell)^{n}\in[\varepsilon,1]$. Finally, Equation 4 holds for $\alpha^{*}(1-\varepsilon)^{2}$ when $q=F(\ell)^{n}\in[0,\varepsilon]$ because the left-hand-side is decreasing in $q$, while the right-hand-side differs by at most the $1-\varepsilon$ factor.
@@
$g(q)~\coloneqq~\int_{\bar{\theta}^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\big\{\bar{\theta}(s),q\big\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~.$
Then, it suffices to show that for any $q\in[0,1]$, it holds
@@
6.2 Verifying the Sufficient Condition
Recall the assumed properties of $\bar{\theta}$ by Equation (1):
@@ 6.2 Verifying the Sufficient Condition
    $\bar{\theta}$ is left-continuous, which implies $\bar{\theta}(\lambda_{1})=1$, and non-increasing; and
@@
    $\int_{z}^{1}\int_{0}^{t}\frac{1}{t}\cdot\bar{\theta}(\max\{s,z\})^{t}\,\mathrm{d}s\,\mathrm{d}t~\geq~\alpha\cdot\bar{\theta}(z)~.$
Consider any $q$ with $\bar{\theta}(\lambda_{2})\leq q\leq\bar{\theta}(\lambda_{1})=1$. Let $z=\bar{\theta}^{-1}(q)$, and then by property 2), we have
$\alpha$ $~\leq~\int_{\bar{\theta}^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\bar{\theta}(\max\{s,\bar{\theta}^{-1}(q)\})^{t}}{\bar{\theta}(\bar{\theta}^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t$
$~=~\int_{\bar{\theta}^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\{\bar{\theta}(s),\bar{\theta}(\bar{\theta}^{-1}(q))\}^{t}}{\bar{\theta}(\bar{\theta}^{-1}(q))}\,\mathrm{d}s\,\mathrm{d}t~.$
By left-continuity of $\bar{\theta}$ and the definition of $\bar{\theta}^{-1}$, we have $\bar{\theta}(\bar{\theta}^{-1}(q))\geq q$. Therefore,
$\alpha~\leq~\int_{\bar{\theta}^{-1}(q)}^{1}\int_{0}^{t}\frac{1}{t}\cdot\frac{\min\{\bar{\theta}(s),q\}^{t}}{q}\,\mathrm{d}s\,\mathrm{d}t~=~g(q)~.$
Finally, we verify $g(q)\geq\alpha$ for $q\in[0,\bar{\theta}(\lambda_{2}))$ by reducing it to the first case.
@@
For any $q\in[0,\bar{\theta}(\lambda_{2})]$, we have
$g(q)\geq g\big(\bar{\theta}(\lambda_{2})\big)~.$
@@
For any $q$ in the range $0,\bar{\theta}(\lambda_{2})]$, we have $\bar{\theta}^{-1}(q)=\lambda_{2}$. Also, observe that $\bar{\theta}(s)\geq\bar{\theta}(\lambda_{2})\geq q$ for $s\in[0,\lambda_{2}]$, and $\bar{\theta}(s)=0$ for $s\in[\lambda_{2},1]$. Hence,
@@
which is decreasing in $q$, implying the desired inequality. ∎
@@ 7 Algorithm for MaxProb: Proof of Theorem 2
This section proves Theorem 2 by establishing the optimal consistency-robustness trade-off for the MaxProb objective. Recall that $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ is the threshold function used by the continuous-time Gilbert-Mosteller algorithm (Example 2), where for each $t\in[0,1]$, the value $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is the unique solution $q$ to
$\int_{t}^{1}\frac{(1-s+sq)^{n-1}-q^{n-1}}{1-s}\,\mathrm{d}s=q^{n-1}~.$ (5)
By the Implicit Sharding technique (Theorem 9), for any $\beta\in[0,\nicefrac{{1}}{{e}}]$, we can focus on the trade-off of the bi-criteria algorithm $\mathcal{A}(\tilde{F},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ in the limiting case when $n\to\infty$, where $\lambda_{1},\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$.
Then, we show that this trade-off is tight within the bi-criteria framework by showing that no algorithm in the form of $\mathcal{A}(\tilde{F},\theta)$ can achieve a better consistency-robustness trade-off.
7.1 Asymptotic Performance Analysis
With Lemma 7, the $\beta$-robustness guarantee of $\mathcal{A}(\tilde{F},\theta_{n}^{\scriptscriptstyle\mathrm{prob}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}})$ follows directly. It remains to show that the algorithm is $\alpha$-consistent as $n\to\infty$, where
To this end, we first derive an explicit expression for the probability that a bi-criteria algorithm accepts the maximum value under a correct prior for finite $n$.
@@ 7.1 Asymptotic Performance Analysis
For any $n\in\mathbb{N}$, any prior $F$, and any threshold function $\theta$, algorithm $\mathcal{A}(\tilde{F}=F,\theta)$ accepts the maximum value with probability
$\Gamma_{n}(\theta)\coloneq\int_{0}^{1}\bigg(\int_{s}^{1}\frac{(1-t+t\theta(s))^{n}-t\theta(s)^{n}}{t(1-t)}\,\mathrm{d}t\;-\;\theta(s)^{n}\bigg)\,\mathrm{d}s\,.$
@@
Changing the order of integration and simplifying yield the expression for $\Gamma_{n}(\theta)$. ∎
With the probability guarantee for general threshold derived in Lemma 14, we next evaluate the consistency achieved by our choice of thresholds for any $\beta\in[0,\nicefrac{{1}}{{\mathrm{e}}}]$.
@@
For any $\beta\in[0,\nicefrac{{1}}{{e}}]$, suppose $\lambda_{1}$, $\lambda_{2}$ are the two roots of $\lambda\ln\lambda=\beta$ with $\lambda_{1}\leq\lambda_{2}$. For any $n\in\mathbb{N}$, recall that $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(t)$ is the unique solution to Equation 5 for any $t\in[0,1]$. Then
$\lim_{n\to\infty}\Gamma_{n}\big(\theta_{n}^{\scriptscriptstyle\mathrm{prob}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)~=~\beta+\int_{\lambda_{1}}^{\lambda_{2}}\int_{s}^{1}\frac{\mathrm{e}^{-\frac{\gamma t}{1-s}}}{t}\,\mathrm{d}t\,\mathrm{d}s~.$
@@
The probability $\Gamma_{n}\big(\theta_{n}^{\scriptscriptstyle\mathrm{prob}}\big|_{\scriptscriptstyle\lambda_{1},\lambda_{2}}\big)$ can be written as
$\int_{0}^{\lambda_{1}}\bigg(\int_{s}^{1}\frac{1}{t}\,\mathrm{d}t-1\bigg)\,\mathrm{d}s+\int_{\lambda_{1}}^{\lambda_{2}}\bigg(\int_{s}^{1}\frac{(1-t+t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s))^{n}-t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)^{n}}{t(1-t)}\,\mathrm{d}t-\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)^{n}\bigg)\,\mathrm{d}s+\int_{\lambda_{2}}^{1}\frac{(1-t)^{n-1}}{t}\,\mathrm{d}s~.$
When $n\to\infty$, the first term equals $\lambda_{1}\ln\lambda_{1}=\beta$ by basic calculus, and the last term tends to 0. It remains to show that, as $n\to\infty$, the inner integrand of the second term converges to
@@
By the definition of $\theta_{n}^{\scriptscriptstyle\mathrm{prob}}$ in Equation 5, the integrand equals
$\quad\int_{s}^{1}\frac{(1-t+t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s))^{n}-t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)^{n}}{t(1-t)}\,\mathrm{d}t\,-\,\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)\int_{s}^{1}\frac{\big(1-t+t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)\big)^{n-1}-\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)^{n-1}}{1-t}\,\mathrm{d}t$
$=\int_{s}^{1}\frac{(1-t+t\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s))^{n-1}}{t}\,\mathrm{d}t~.$
Finally, we use the asymptotic expansion of the threshold function proven in Appendix C:
@@
$\theta_{n}^{\scriptscriptstyle\mathrm{prob}}(s)~=~\frac{1}{1+\frac{\gamma}{(n-1)(1-s)}}+o\!\left(\frac{1}{n}\right)$